OpenStudy (anonymous):

WILL BECOME A FAN AND GIVE A MEDAL What is the largest rectangle with lower base on the x-axis and upper vertices on the curve y=12-x^2? (A) 8 (B) 12 (C) 16 (D) 32 (E) 48 Could you please show your work as well as select an answer, thank you so much!

OpenStudy (anonymous):

First, let's state what we know. We know the general area of a rectangle and the curve we are given, so let's put them down: y = 12 - x^2 Area = xy Now, substitute in what we're given from the curve into the Area formula: Area = x(12 - x^2) = -x^3 + 12x To find the max area, we should differentiate and set it equal to zero to solve for our critical value(s). Area ' = -3x^2 + 12 Area ' = 0 = -3x^2 + 12 -12 = -3x^2 4 = x^2 x = +/- 2 Now, we want to maximize area! This means that the x-value that we want should generate the larger of the areas when substituted into the area expression we defined earlier: Area = -x^3 + 12x *if x = -2* Area = -(-2)^3 + 12(-2) = -(-8) - 24 = 8 - 24 = -16 <-- well, we know that isn't possible to have a negative area!! x = 2 must be our solution, let's plug it in: Area = -x^3 + 12x *if x = 2* Area = -(2)^3 + 12(2) = -(8) + 24 = 16 So our solution is that the maximum area of the rectangle under this curve is 16 sq. units. That corresponds to answer C.

OpenStudy (anonymous):

oh okay thank you so much!